izukumidoryia:

A monument in a park is shaped like a square pyramid. The dimensions are shown in the net. What is the surface area of the monument, including the base? A. 4 m² B. 12 m² C. 14 m² D. 24 m² Can you help my bro?

izukumidoryia:

1 attachment
iiMaddii:

What answer r ya'll thinking >?

izukumidoryia:

where sayin D

iiMaddii:

But I was thinking a different answer

iiMaddii:

Here's Some explanation why .

Shuki:

arent you supposed to do 2.5 x 3.14/ 2

Shuki:

for 1 triangle? and then add that answer 4 more times then add 4?

iiMaddii:

hm

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shuki arent you supposed to do 2.5 x 3.14/ 2 \(\color{#0cbb34}{\text{End of Quote}}\) what? \( \pi \) is not necessary at all -- do you know what the formula to find the area of a triangle is?

Shuki:

I may be confusing it with the eqauion factors I'm doing now

izukumidoryia:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken \(\color{#0cbb34}{\text{Originally Posted by}}\) @Shuki arent you supposed to do 2.5 x 3.14/ 2 \(\color{#0cbb34}{\text{End of Quote}}\) what? \( \pi \) is not necessary at all -- do you know what the formula to find the area of a triangle is? \(\color{#0cbb34}{\text{End of Quote}}\) A=12bh where b is the base and h is the height of the triangle.

Florisalreadytaken:

perhaps -- so this would be the formula \[ \color{dodgerblue}{\triangle}\color{lightskyblue}{A} \color{lightskyblue}{\scriptsize\text{rea}} =\frac{\color{chocolate}{b} \times \color{teal}{h}} {2} \] \(\color{chocolate}{b} =2\) \( \color{teal}{h}=2.5 \)

izukumidoryia:

he got it

Florisalreadytaken:

sweet as!